The function f is defined by f(x)= (5+x)/(-2+3x). Find f(4x)
Answer by amarjeeth123(570)

f(x)= (5+x)/(-2+3x)
f(4x)=(5+4x)/(-2+3(4x))=(5+4x)/(-2+12x)